Nova Company’s total overhead cost at various levels of activity are presented below: Month Machine- Hours Total Overhead Cost April 47,000 $ 180,200 May 37,000 $ 156,200 June 57,000 $ 204,200 July 67,000 $ 228,200 Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 37,000 machine-hour level of activity is: Utilities (variable) $ 48,100 Supervisory salaries (fixed) 55,000 Maintenance (mixed) 53,100 Total overhead cost $ 156,200 Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements. Required: 1. Estimate how much of the $228,200 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $228,200 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) 2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. 3. Express the company’s total overhead cost in the form Y = a + bX. 4. What total overhead cost would you expect to be incurred at an activity level of 42,000 machine-hours?
Answer #1
Month Machine Hours Overhead Cost
April               47,000 $180,200
May               37,000 $156,200
June               57,000 $204,200
July               67,000 $228,200
37000 Variable cost(Utilities) $48,100
Fixed cost(supervisor salary) $55,000
Maint. Cost(mixed) $53,100
Total overhead $156,200
Requirement-1 Maint. Cost in July = 228,200
Machine hours = 67,000
Utilities cost(variable) $87,100 (48,100/37,000*67,000)
Fixed Cost(supervisor salary) $55,000
$142,100
hence, remaining cost is maintenance cost $86,100 (228,200-142,100)
Requirement -2 Variable cost per unit as per high low method = (Highest Activity Cost - Lowest Activity Cost)/(Highest activity units - Lowest activity units)
(228,200-156,200)/(67,000-37,000)
Variable cost per unit as per high low method = 2.4
Fixed cost = Highest activity cost – (Variable cost per unit x Highest activity units)
Fixed cost = 228,200 - (2.4*67000)
Fixed cost = 228,200 - 160,800
Fixed cost = $ 67,400
Y = 67,400 + 2.40*X
Where Y = total cost
X = Activity hours
Requirement -3 Alternative approach Differencial Variable portion at 37,000 level Fixed portion
Activity level -> 37000 67000 30000
utility cost $48,100 $87,100 39000 1.3 48100 0
(87,100-48,100) (39,000/30,000) (1.3*37,000)
maint $53,100 $86,100 33000 1.1 40700 $12,400
(86,100-53,100) (33,000/30,000) (1.1*37,000) (53100-40,700)
Hence total fixed overhead = 55,000+12,400
Hence total fixed overhead = 55,000+12,400
Hence total fixed overhead = 67,400
Y = 67,400 + 2.40*X
Where Y = total cost
X = Activity hours
Requirement -4 Y = 67,400 + 2.40*X
Y = 67,400 + 2.40*42,000
Y = 67,400 + 100,800
Y = $ 168,200
Hence, total cost = 168,200